# Hunt Globally: Wide Search AI Agents for Drug Asset Scouting in Investing, Business Development, and Competitive Intelligence

### v2

@@ Abstract
Bio-pharmaceutical innovation has shifted: many new drug assets now originate outside the United States and are disclosed primarily via regional, non-English channels. Recent data suggests over 85% of patent filings originate outside the U.S., with China accounting for nearly half of the global total; a growing share of scholarly output is also non-U.S. Industry estimates put China at $\sim 30\%$ of global drug development, spanning 1,200+ novel candidates. In this high-stakes environment, failing to surface "under-the-radar" assets creates multi-billion-dollar risk for investors and business development teams, making asset scouting a coverage-critical competition where speed and completeness drive value. Yet today’s Deep Research AI agents still lag human experts in achieving high-recall discovery across heterogeneous, multilingual sources without hallucinations. We propose a benchmarking methodology for drug asset scouting and a tuned, tree-based self-learning Bioptic Agent aimed at complete, non-hallucinated scouting. We construct a challenging completeness benchmark using a multilingual multi-agent pipeline: complex user queries paired with ground-truth assets that are largely outside U.S.-centric radar. To reflect real deal complexity, we collected screening queries from expert investors, BD, and VC professionals, and used them as priors to conditionally generate benchmark queries. For grading, we use LLM-as-judge evaluation calibrated to expert opinions. On this benchmark, our Bioptic Agent achieves 79.7% F1 score, outperforming Claude Opus 4.6 (56.2%), Gemini 3 Pro + Deep Research (50.6%), OpenAI GPT-5.2 Pro (46.6%), Perplexity Deep Research (44.2%), and Exa Websets (26.9%). Performance improves steeply with additional compute, supporting the view that more compute yields better results.
@@ Introduction
This paper expands our previous work on benchmarking LLM-based agents for competitive landscape mapping in drug asset due diligence [Vinogradova et al., 2025] by moving from competitor discovery within a single diligence context to open-world, multilingual, “find-all” asset scouting with controlled query intent and difficulty. In a head-to-head evaluation against state-of-the-art commercial deep-research baselines, Bioptic Agent achieves 79.7% F1-score (harmonic mean of precision and recall; higher is better) and substantially outperforms Claude Opus 4.6 high at 56.2%, Gemini 3 Pro with Deep Research at 50.6%, OpenAI GPT-5.2 Pro high at 46.6%, Perplexity Deep Research at 44.2%, and Exa Websets at 26.9%, supporting the conclusion that investor-grade and BD-grade scouting requires completeness-oriented search control, lossless candidate tracking, and expert-aligned validation rather than increased browsing compute or better synthesis alone.
@@
    Deep research agents: systems optimized for prolonged browsing + synthesis (e.g., Gemini 3 Pro Deep Research, Perplexity Deep Research).
@@
Table 2 shows that Bioptic Agent achieves the best overall performance, with an F1-score of $0.797$, substantially outperforming all tested state-of-the-art search agents. The strongest baseline is Claude Opus 4.6 with an F1-score of $0.562$, followed by Gemini 3 Pro Deep Research with an F1-score of $0.506$, o4-mini-deep-research with $0.489$, GPT-5.2 Pro high with $0.466$, Perplexity Sonar Deep Research high with $0.442$, GPT-5.2 high with $0.287$, and Exa Websets with $0.269$. Bioptic Agent’s advantage is driven by simultaneously high precision of $0.877$ and high recall of $0.730$, whereas competing systems exhibit markedly lower recall and more limiting precision–recall tradeoff.
@@
We include OpenAI Deep Research (o4-mini) in the sequential scaffold ablation to test a natural hypothesis suggested by Table 2: since Gemini 3 Pro Deep Research is the second strongest after Bioptic Agent, perhaps simply running a state-of-the-art Deep Research agent longer—by iterating a sequential “find more” loop—could close the gap to Bioptic Agent. Directly applying many sequential epochs to Gemini 3 Pro Deep Research is costly, so we use o4-mini-deep-research as a cost-efficient alternative that is comparable in terms of F1-score and itself a Deep Research scaffolded agent. Empirically, however, the sequentially wrapped o4-mini-deep-research still improves more slowly and saturates earlier than Bioptic Agent, similarly to how the gpt-5.2-based scaffold is saturated, consistent with diminishing returns once the most obvious sources and search angles are exhausted. This demonstrates that the gains are not in simply running a state-of-the-art general agent longer; rather, Bioptic Agent’s advantage comes from its specialized tree-based exploration, self-reflection, and self-learning scaffolding beyond naive sequential iteration.
@@ Conclusion
On this benchmark, Bioptic Agent achieves the best overall performance with F1-score of 79.7% and substantially outperforms state-of-the-art commercial deep-research baselines (Table 2), including Claude Opus 4.6 high (56.2%), Gemini 3 Pro Deep Research (50.6%), o4-mini-deep-research (48.9%), GPT-5.2 Pro high (46.6%), Perplexity Sonar Deep Research high (44.2%), and Exa Websets (26.9%). Figure 1 shows a clear quality–time tradeoff: Bioptic Agent improves rapidly early and then approaches a plateau near $\sim 0.80$ F1, while sequential “run longer” scaffolds improve more slowly and plateau earlier at lower quality.

### v3

@@ Abstract
Bio-pharmaceutical innovation has shifted: many new drug assets now originate outside the United States and are disclosed primarily via regional, non-English channels. Recent data suggests over 85% of patent filings originate outside the U.S., with China accounting for nearly half of the global total; a growing share of scholarly output is also non-U.S. Industry estimates put China at $\sim 30\%$ of global drug development, spanning 1,200+ novel candidates. In this high-stakes environment, failing to surface "under-the-radar" assets creates multi-billion-dollar risk for investors and business development teams, making asset scouting a coverage-critical competition where speed and completeness drive value. Yet today’s Deep Research AI agents still lag human experts in achieving high-recall discovery across heterogeneous, multilingual sources without hallucinations. We propose a benchmarking methodology for drug asset scouting and a tuned, tree-based self-learning Bioptic Agent aimed at complete, non-hallucinated scouting. We construct a challenging completeness benchmark using a multilingual multi-agent pipeline: complex user queries paired with ground-truth assets that are largely outside U.S.-centric radar. To reflect real deal complexity, we collected screening queries from expert investors, BD, and VC professionals, and used them as priors to conditionally generate benchmark queries. For grading, we use LLM-as-judge evaluation calibrated to expert opinions. On this benchmark, our Bioptic Agent achieves 79.7% F1 score, outperforming Claude Opus 4.6 (56.2%), Gemini 3.1 Deep Think (59.2%), Gemini 3.1 Pro Deep Research (58.6%), OpenAI GPT-5.2 Pro (46.6%), Perplexity Deep Research (44.2%), and Exa Websets (26.9%). Performance improves steeply with additional compute, supporting the view that more compute yields better results.
@@ Introduction
This paper expands our previous work on benchmarking LLM-based agents for competitive landscape mapping in drug asset due diligence [Vinogradova et al., 2025] by moving from competitor discovery within a single diligence context to open-world, multilingual, “find-all” asset scouting with controlled query intent and difficulty. In a head-to-head evaluation against state-of-the-art commercial deep-research baselines, Bioptic Agent achieves 79.7% F1-score (harmonic mean of precision and recall; higher is better) and substantially outperforms Claude Opus 4.6 high at 56.2%, Gemini 3.1 Deep Think at 59.2%, Gemini 3.1 Pro Deep Research at 58.6%, OpenAI GPT-5.2 Pro high at 46.6%, Perplexity Deep Research at 44.2%, and Exa Websets at 26.9%, supporting the conclusion that investor-grade and BD-grade scouting requires completeness-oriented search control, lossless candidate tracking, and expert-aligned validation rather than increased browsing compute or better synthesis alone.
@@
    Deep research agents: systems optimized for prolonged browsing + synthesis (e.g., Gemini 3.1 Pro Deep Research, Gemini 3 Pro Deep Research, Perplexity Deep Research).
@@
    Scientific Research & Reasoning Agents: Gemini 3.1 Deep Think evaluated through their publicly available UI app.
  6. 6.
@@
Gemini 3.1 Deep Think $0.636$ $0.554$ $0.592$
Gemini 3.1 Pro Deep Research $0.545$ $0.634$ $0.586$
@@
Table 2 shows that Bioptic Agent achieves the best overall performance, with an F1-score of $0.797$, substantially outperforming all tested state-of-the-art search agents. The strongest baseline is Gemini 3.1 Deep Think with an F1-score of $0.592$, followed by Gemini 3.1 Pro Deep Research with an F1-score of $0.586$, Claude Opus 4.6 with $0.562$, Gemini 3 Pro Deep Research with $0.506$, o4-mini-deep-research with $0.489$, GPT-5.2 Pro high with $0.466$, Perplexity Sonar Deep Research high with $0.442$, GPT-5.2 high with $0.287$, and Exa Websets with $0.269$. Bioptic Agent’s advantage is driven by simultaneously high precision of $0.877$ and high recall of $0.730$, whereas competing systems exhibit markedly lower recall and more limiting precision–recall tradeoff.
@@
We include OpenAI Deep Research (o4-mini) in the sequential scaffold ablation to test a natural hypothesis suggested by Table 2: since Gemini 3.1 Pro Deep Research is the second strongest after Bioptic Agent, perhaps simply running a state-of-the-art Deep Research agent longer—by iterating a sequential “find more” loop—could close the gap to Bioptic Agent. Directly applying many sequential epochs to Gemini 3 Pro Deep Research is costly, so we use o4-mini-deep-research as a cost-efficient alternative that is comparable in terms of F1-score and itself a Deep Research scaffolded agent. Empirically, however, the sequentially wrapped o4-mini-deep-research still improves more slowly and saturates earlier than Bioptic Agent, similarly to how the gpt-5.2-based scaffold is saturated, consistent with diminishing returns once the most obvious sources and search angles are exhausted. This demonstrates that the gains are not in simply running a state-of-the-art general agent longer; rather, Bioptic Agent’s advantage comes from its specialized tree-based exploration, self-reflection, and self-learning scaffolding beyond naive sequential iteration.
@@ Conclusion
On this benchmark, Bioptic Agent achieves the best overall performance with F1-score of 79.7% and substantially outperforms state-of-the-art commercial deep-research baselines (Table 2), including Claude Opus 4.6 high (56.2%), Gemini 3.1 Deep Think (59.2%), Gemini 3.1 Pro Deep Research (58.6%), Gemini 3 Pro Deep Research (50.6%), o4-mini-deep-research (48.9%), GPT-5.2 Pro high (46.6%), Perplexity Sonar Deep Research high (44.2%), and Exa Websets (26.9%). Figure 1 shows a clear quality–time tradeoff: Bioptic Agent improves rapidly early and then approaches a plateau near $\sim 0.80$ F1, while sequential “run longer” scaffolds improve more slowly and plateau earlier at lower quality.